In 2015, the Washington Nationals baseball team signed pitcher Max Scherzer to a contract to
play for the team for 7 years. He would be paid $15 million dollars per year for 14 years—an ad- ditional 7 years beyond the end of the time he would be committed to play for the Nationals. The contract was widely reported as being worth $210 million (or $15 million per year * 14 years). One baseball writer argued, though, that “this deal serves as a nice reminder that the payment terms of a deal can have an impact on the actual value of the contract.”

a. What does the writer mean by the “actual value of the contract”? Isn’t $210 million the actual value of the contract? Briefly explain.

b. the writer notes :“For alot of reasons,money today is worth more than money in the future, and the further in the future you go, the less money is worth.” What are the reasons that money today is worth more than money in the future? Briefly explain how we calculate the value today of money we will receive in the future.

c. Assume for simplicity that Scherzer receives his salary of $15 million per year for 2018, 2019, and 2020 at the end of each calendar year. At the beginning of 2018, what is the present value of the salaries he will receive for these three years? Use an interest rate of 7% in your calculation.
Answer #1

a. The value of the contract is measured on a particular day. Due to the concept of the time value of money, The $15 million received for this year is worth much more that the $15 million received at the end of the 14th year, as it can be invested in various avenues and interest earned from it.

Also, the actual value of an amount which is received in future is lesser today. hence all the future payments have to be discounted at the rate of interest to get the present and hence the actual value of the contract.

b. Money today is worth more than money in the future as it can be invested in various avenues and interest earned from it. The way to calculated the value of a sum of money which we would get in the future is to look at the interest that we can get if we invest money in some avenue - a bank, a fund, etc.. lets say it is 7%. Then the value today is found by discounting the future value by the interest rate.

Example, if the amount received after nth year = 100, then present value = 100/(1+r)n   

where r is the rate of interest, n is the number of years

c. Using the formula in part b, we can calculate the the present value of the salaries he will receive for these three years as

= 15/(1+0.07)1 + 15/((1+0.07)2 + 15/((1+0.07)3

= 14.019 + 13.1016 + 12.244

=$39.29 million